if a student were paid .25 for each math test taken, and was fined .50 for each failed math test; and he passed 7x's as many tests as he failed. How many tests did he fail, if he was paid $3.75?
Answer by stanbon(75887)

if a student were paid .25 for each math test taken, and was fined .50 for each failed math test; and he passed 7x's as many tests as he failed. How many tests did he fail, if he was paid $3.75?
-----
p = 7f
0.25p - 0.50f = 3.75
---------------------------------
Modify::
p = 7f
p = 2f + 15
----
7f = 2f+15
5f = 15
# failed = 3
----
# passed = 7*3 = 21
--------------
Cheers,
Stan H.
--------------